OpenStudy (anonymous):

Solve by factoring x^2-x-90=0

OpenStudy (anonymous):

-10 and 9

OpenStudy (anonymous):

(x-10)(x+9)

OpenStudy (anonymous):

set both equal to 0 and solve for x x = 10 and -9

OpenStudy (accessdenied):

So, you're looking for factors of -90 that add up to -1. If the sum of two numbers is negative, then the larger one would be negative. Try listing the factors of -90 and adding them together. Then you may rewrite the equation as (x - a)(x - b) = 0 you can then do x-a = 0, x-b = 0 and solve for the x's

OpenStudy (anonymous):

Does that help? Rather you should be practicing how to get to that...it gets easier